TE1AES004 cavity history Updated on:  1/12/2010
     
Date Activity Notes
  Order Link to Vendor data
  QA Link to ECS
  Fabrication
  QA
  BCP 107 micron BCP procedure by Holly.jpg
  left in DI water
11/28/2007 HPR 2 hours By D. Meidlinger, Z. Conway and Y. Xie
12/1/2007 Dry over night Cavity prepared by D. Meidlinger, Z. Conway and Y. Xie
12/3/2007 RF test 1 Tested by G. Wu, N. Dhanaraj, Y. Xie and D. Meidlinger
EP 65 micron at ANL Kelly and Gerbick
HPR at A0  
Assemble at MP9  
IB1 pump down  
IB1 Q-T measurement Q-T
9/30/2008 IB1 Q-Eacc measurement 33 MV/m limited by FE and Q-slope
10/3/2008 IB1 RF test 2 summary Q-Eacc Ozlis and Wu
Pumping line vacuum leak, up to torr at one time at IB1
10/15/2008 IB1 48h 120C baking intermittently applied in 3 days
10/23/2008 IB1 RF test 3 23 MV/m limited by heavy field emission Ozelis
11/20/2008 disassembled and HPR at A0
12/8/2008 IB1 RF Test 4 33 MV/m limited by FE and Q-slope Ozelis
1/15/2009 cavity imaging FNAL_ICB Big pit but no quench (click to see) M. Ge and G. Wu
2/4/2009 HPR 2 cycles 170 minutes ?   Bice and Olis
2/5/2009 Assembly and pumpdown quite slow pumpdown, limited <1.5 torr Bice and Wu
2/6/2009 Leak check, T-map, in dewar   Bice, Wu, Carter, Mukherjee, Massengill
2/9/2009 RF test 5 39 MV/m, quench limited, slight field emission processed away Ozelis and Wu
3/13/2009 molding test  
4/6/2009 US soap, ethanol rinsing, HPR
1st Assembly, 2nd HPR
4/7/2009 Final assembly, pumpdown
4/10/2009 RF test 6 to study molding material 39.3 MV/m limited by quench, same as previous test. No field emission
9/18/2009 RF test 7 to investigate the A0 RF system Field emission appeared, Q/Eacc dropped a little Ge and Khabibouline  
10/2/2009 RF test 8 to locate quench location Ge and Harms
11/6/2009 RF test 9 to repeat T-map Ge and Harms
12/23/2009 RF test 10 to test second sound 33.8 MV/m
1/11/2010 RF test 11 to repeat second sound test 37 MV/m

Legal Notices